Write an algebraic equation for the following:
1.Donna has c dimes and t quarters in her bank. How much money in cents does she have?
2.If a gallons of gas cost m dollars, what is the price per gallon?
3.If c represents an odd number, what is the next larger odd number?
4.Willie is y years old, and his father is 4 years less than four times Willie's age. What is the sum of their ages?

Answer by Theo(13342)

1.Donna has c dimes and t quarters in her bank. How much money in cents does she have?
-----
she will have 10*c + 25*q cents.
example:
she has 3 dimes and 1 quarter.
This equals 10*3 + 25*1 cents which equals 30 + 25 cents which equals 55 cents.
-----
2.If a gallons of gas cost m dollars, what is the price per gallon?
-----
The price per gallon is the cost divided by the number of gallons.
price per gallon equals m / a.
example:
cost for gas is $10.00
number of gallons bought is 5
cost per gallon is $10.00/5 = $2.00 per gallon.,
-----
3.If c represents an odd number, what is the next larger odd number?
-----
Next odd number would be c + 2
example:
c = 5, 5+2 = 7, 7+2 = 9, etc.
-----
4.Willie is y years old, and his father is 4 years less than four times Willie's age. What is the sum of their ages?
-----
willie is y years old
His father is 4*y - 4
sum of their ages is 6 + 4y -4 = 5y - 4
example:
willie is 10 years old.
his father is 36 years old.
sum of their ages is 46 years.
y = 10
4*y-4 = 40-4 = 36
5y-4 = 46
-----